# Live-SWE-agent: Can Software Engineering Agents Self-Evolve on the Fly?

### v2

@@ Abstract
Figure 1: SWE-bench Verified and SWE-Bench Pro single attempt (w/o test-time scaling) results.
@@ 1 Introduction
Large Language Models (LLMs) have rapidly progressed from simple code autocompletion austin2021program; chen2021codex; li2022competition to interactive agents capable of navigating repositories, running tests, and submitting patches end‑to‑end yang2024sweagent; wang2024openhands; zhang2024autocoderover; xia2024agentless; carbonneaux2025cwm; liu2024large. Early conversational repair systems leveraged environment feedback to iteratively refine candidate fixes xia2024automated, while subsequent agentic frameworks (such as SWE-agent yang2024sweagent and OpenHands wang2024openhands) augmented LLMs with terminals, editors, and search, enabling multi‑step tool use on complex repositories. At the same time, complementary “more agentless” pipelines (e.g., Moatless moatless and Agentless xia2024agentless) argue that much of the perceived complexity in agent scaffold design can be replaced by dedicated prompting and workflow design.
@@
  • Live, on-the-fly self-evolution. We present Live-SWE-agent, the first live software agent that can autonomously self-evolve its own scaffold implementation on the fly when solving real-world issues, without any offline training or extra pipelines.
  • Minimal and general implementation. Our current implementation adopts a minimal and general design, where the agent starts from a minimal bash-only scaffold (e.g., mini-SWE-agent) and self-improves on the fly by creating general or customized tools. The design is compatible with any existing software agent loop or LLM with negligible overhead.
  • State-of-the-art open-source agent. On SWE-bench Verified and SWE-Bench Pro, Live-SWE-agent reaches 77.4% and 45.8% solve rates respectively, outperforming leading open-source agents and commercial systems, without any test-time scaling. To our knowledge, our SWE-Bench Pro result is the best reported to date.
  • Comprehensive analysis. We include detailed studies of when and why on-the-fly tool creation helps, how it improves efficiency, and how it compares to other designs like fixed-tool agents, workflow-based systems, and offline self-improving methods. Notably, compared with existing self-improving agents, we achieve much better issue-solving performance (65.0% issue resolution rate on the SWE-bench Verified-60 subset vs. DGM’s 53.3%) while being significantly less costly (zero offline cost).
@@ 3 Experimental Setup
Implementation. While Live-SWE-agent is general across different agent scaffolds, we implement Live-SWE-agent on top of the popular mini-SWE-agent (yang2024sweagent) framework. As such, we retain the hyperparameters used in mini-SWE-agent by default (i.e., temperature 0.0, maximum step limit of 250, and maximum cost of \$3 per issue). We choose mini-SWE-agent to build on since it is not only simplistic (with $\sim$100 lines of code and only accessing bash commands) but also widely used. Unless otherwise stated, we use Claude 4.5 Sonnet (claude-sonnet-4-5-20250929(sonnet4.5) in our experiments and sample one patch per issue.
@@ 4.1 Main Results
SWE-bench Verified. Table 1 shows the results of Live-SWE-agent and prior agent approaches on SWE-bench Verified. We first observe that compared with mini-SWE-agent, across three different LLM backends, Live-SWE-agent consistently achieves a higher resolve rate. This demonstrates the improvement in performance enabled by allowing agents to create and use their own custom tools on the fly. Furthermore, Live-SWE-agent is able to achieve this with only a minimal increase in cost compared with the base mini-SWE-agent. In certain cases (e.g., in GPT-5), we even observed slight cost savings where the agent can improve the solving efficiency by replacing complex, multi-turn commands with customs tools that achieve the same functionality.
We also demonstrate the comparison between Live-SWE-agent and the best-performing agentic tools, including both state-of-the-art open-source solutions and proprietary commercial scaffolds, on SWE-bench Verified (Figure 1). We report both the base agent performance and, when applicable, the result with test-time scaling. We observe that Live-SWE-agent using Gemini 3 Pro achieves a solve rate of 77.4% without test-time scaling, outperforming all existing agents on the SWE-bench Verified leaderboard [^1] https://www.swebench.com including state-of-the-art commercial solutions at the time of writing.
@@
SWE-Bench Pro. We also evaluate Live-SWE-agent on the public set of SWE-Bench Pro, containing 731 problems across 11 repositories and four programming languages (Python, Go, TypeScript, and JavaScript). Table 3 shows the performance of Live-SWE-agent compared to the state-of-the-art baseline of SWE-agent on SWE-Bench Pro. Different from mini-SWE-agent, which only has access to basic bash commands, SWE-agent provides handcrafted file viewing and editing tools and is the best performing approach on the SWE-Bench Pro leaderboard [^2] https://scale.com/leaderboard/swe_bench_pro_public. We observe that Live-SWE-agent is able to achieve better performance compared with SWE-agent, a specially designed agent implemented with close to 7,000 lines of code. We also compare with all other top-performing agents and LLMs in Figure 1 and observe that Live-SWE-agent is able to achieve the new state-of-the-art performance of 45.8% resolve rate on SWE-Bench Pro. This demonstrates that Live-SWE-agent can lead to improved results compared with techniques that start with a manually-crafted set of tools by enabling the agent to build and use custom tools on the fly to perform self-evolution.
@@ 4.2 Tools Analysis
In addition to looking at the types of tools created, we also examine how the tools created can change across different repositories and programming languages. Figure 4(b) shows the visualization of the tools created in SWE-Bench Pro categorized based on repositories. In SWE-Bench Pro, there are 11 different repositories, and not surprisingly, there are common tools created across different repositories (i.e., the overlapping clusters in the figure). On the other hand, there are also specific tools for certain repositories, for example we see a distinct cluster (highlighted in orange diamonds) for openlibrary, a codebase for cataloging literature. Since openlibrary deals with lots of raw data of a specialized format, the custom tools generated by Live-SWE-agent are specifically tailored for that. We also observe similar results when we break down the tools based on programming language of the repositories in Figure 4(c). Different problems, environments, and languages may require completely different specialized tools, highlighting again how using the same set of tools for all problems is suboptimal.
@@ 4.3 Ablation
Different LLM backends. We now examine the effect of using different LLM backends for Live-SWE-agent. Table 5 compares the performance of the base mini-SWE-agent and Live-SWE-agent as we vary the LLM used on the same 50 problems used in the previous ablation experiment. We first observe that there are weaker LLMs for which using Live-SWE-agent to create custom tools on the fly decreases performance. In particular, we see that for GPT-5-Nano, the results are even significantly worse when using Live-SWE-agent compared to base mini-SWE-agent. After examining the trajectories, we found that GPT-5-Nano fails to understand the goal of creating custom tools and is often stuck in a loop thus leading to reduction in performance. This indicates that weaker LLMs may lack the high-level reasoning capabilities to synthesize useful tools on the fly. However, we see that Live-SWE-agent is able to improve more over the base mini-SWE-agent as we start to use more powerful models. Specifically, state-of-the-art LLMs such as Claude 4 Sonnet and GPT-5 achieves the highest relative resolve rate improvement when using Live-SWE-agent compared with mini-SWE-agent. This demonstrates the generalizability of Live-SWE-agent across high-performing LLMs and the potential for Live-SWE-agent to further improve performance especially as we continue to build more and more powerful LLMs.
@@ 4.4 Discussion and Future Work
Towards general, on the fly self-evolution. In this work, Live-SWE-agent primarily focuses on enabling agents to self-evolve through custom tool creation and usage. As discussed previously, the key idea of Live-SWE-agent, to improve and modify the agent itself on the fly, extends beyond creating new tools to also modifying the entire agent implementation. This includes the overall system prompt of the agent, how the agent interacts with the environment, and even the concrete workflow of the agent when it attempts to solve the issue. We believe that the entire agentic loop, much like any other software, can be modified and improved especially on the fly based on feedback and insights gained from solving a problem. Furthermore, we hope to extend the self-evolution loop across different tasks as well. Instead of discarding each evolved agent after a task is completed, we can save and serialize the useful tools and insights (via concepts like Skills skills) for future tasks. The agent can then load these useful tools and insights obtained while solving previous tasks on the fly to improve performance and support continuous self-evolution across tasks.
@@ 5.1 Software Engineering Agents
Due to the huge design space for software agents, it can be extremely challenging and costly to build an optimal agent scaffold. As a result, a number of self-improving software agents have been proposed very recently, including the Self-Improving Coding Agent (SICA) (robeyns2025sica), Darwin-Gödel Machine (DGM) (zhang2025darwin), and Huxley-Gödel Machine (HGM) (wang2025huxley). However, such self-improving agents require costly offline training on known benchmarks, and may not be generalize well across different LLMs, benchmarks, and issue types. Beyond the software engineering domain, prior work llmastool; wang2024voyager; qiu2025alitageneralistagentenabling; qian2024creatortoolcreationdisentangling; wang2024troveinducingverifiableefficient explored using LLMs to create tools for general reasoning or embodied tasks, but they do not target real-world software engineering problems. In contrast, in this paper, we propose Live-SWE-agent, the first software agent capable of performing practical self-evolution on-the-fly during runtime when solving real-world issues. In this way, Live-SWE-agent requires no offline training at all and can be easily generalized to different LLMs and issue domains. Moreover, it also demonstrated superior performance compared to existing self-improving software agents.
@@ 5.2 Benchmarks for Software Engineering Agents
To evaluate and demonstrate the performance of software engineering agents, a large number of datasets have been proposed. SWE-bench (jimenez2023swe) is one of the earliest and most widely used benchmark datasets for software agents. Besides the initial SWE-bench dataset, which includes thousands of real-world GitHub issues, researchers have also built curated subsets with high-quality and representative issues to support faster and more reliable evaluation, including SWE-bench Lite (jimenez2023swe) and SWE-bench Verified (sbv). Since SWE-bench mostly focused on Python projects, researchers have proposed a number of SWE-bench-style benchmarks for projects in multiple languages for more comprehensive agent evaluation, including SWE-PolyBench (rashid2025swe), SWE-bench Multilingual (sbm), and Multi-SWE-bench (zan2025multi). Also, SWE-bench relies heavily on manual effort for collecting benchmark instances and setting up executable environments, and is hardly scalable; as a result, researchers have also leveraged software agents to streamline issue collection and environment setup to build live and scalable benchmarks, including SWE-bench-Live (zhang2025swe). More recently, Scale AI has also constructed SWE-Bench Pro (deng2025swe), which aims to capture more realistic, complex, enterprise-level issues than SWE-bench. For rigorous evaluation of Live-SWE-agent, our study involved multiple widely used benchmarks, including SWE-bench Verified, SWE-bench Multilingual, and SWE-bench Pro.
@@ Appendix A Additional experimental settings
We now describe additional experimental settings used by Live-SWE-agent for evaluation. As discussed in Section 3, we implement Live-SWE-agent on top of the mini-SWE-agent framework. We use the same default hyperparameter of mini-SWE-agent at the time of writing (i.e., temperature 0.0, maximum step limit of 250, and maximum cost of \$3 per issue).
For Gemini 3 Pro, we use temperature of 1 since it is recommended by the developers [^4] https://ai.google.dev/gemini-api/docs/gemini-3. For any OpenAI models in our experiments (e.g., GPT-5, GPT-5-Mini, GPT-5-Nano), since temperature 0.0 sampling is not supported, we use temperature of 1. To the best of our knowledge, this is also the case used in mini-SWE-agent when they use OpenAI models.

### v3

@@ Abstract
Figure 1: SWE-bench Verified and SWE-Bench Pro results (single attempt w/o test-time scaling)
@@ 1 Introduction
Large Language Models (LLMs) have rapidly progressed from simple code autocompletion austin2021program; chen2021codex; li2022competition; xia2022less to interactive agents capable of navigating repositories, running tests, and submitting patches end‑to‑end yang2024sweagent; wang2024openhands; zhang2024autocoderover; xia2024agentless; carbonneaux2025cwm; liu2024large. Early conversational repair systems leveraged environment feedback to iteratively refine candidate fixes xia2024automated, while subsequent agentic frameworks (such as SWE-agent yang2024sweagent and OpenHands wang2024openhands) augmented LLMs with terminals, editors, and search, enabling multi‑step tool use on complex repositories. At the same time, complementary “more agentless” pipelines (e.g., Moatless moatless and Agentless xia2024agentless) argue that much of the perceived complexity in agent scaffold design can be replaced by dedicated prompting and workflow design.
@@
  • The first live software agent. We present Live-SWE-agent, the first live software agent that can autonomously self-evolve its own scaffold implementation on the fly when solving real-world issues, without any offline training or extra pipelines.
  • Minimal and general implementation. Our current implementation adopts a minimal and general design, where the agent starts from a minimal bash-only scaffold (e.g., mini-SWE-agent) and self-improves on the fly by creating general or customized tools. The design is compatible with any existing software agent loop or LLM with negligible overhead, has been publicly available at: https://github.com/OpenAutoCoder/live-swe-agent.
  • State-of-the-art performance. On SWE-bench Verified and SWE-Bench Pro, Live-SWE-agent reaches 77.4% and 45.8% solve rates (without any test-time scaling), respectively, outperforming all existing open-source agents and commercial systems at the time of writing. To our knowledge, our SWE-Bench Pro result is also the best reported to date.
  • Comprehensive analysis. We include detailed studies of when and why on-the-fly tool creation helps, how it improves efficiency, and how it compares to other designs like fixed-tool agents, workflow-based systems, and offline self-improving methods. Notably, compared to existing self-improving agents, we achieve much better performance (65.0% solve rate on the SWE-bench Verified-60 subset vs. DGM’s 53.3%) while being significantly less costly (no offline cost).
  • Unified leaderboard. For software tasks, recent LLMs are often benchmarked using manually engineered, proprietary agent scaffolds, making fair model comparison hard. Live-SWE-agent offers an open, unified, and powerful scaffold that enables genuinely fair, apples-to-apples comparison for future model releases. We are maintaining a leaderboard of recent models evaluated on real-world software tasks using Live-SWE-agent at: http://live-swe-agent.github.io
@@ 3 Experimental Setup
Implementation. While Live-SWE-agent is general across different software agent scaffolds, we implement Live-SWE-agent on top of the popular mini-SWE-agent (yang2024sweagent) framework. As such, we retain the hyperparameters used in mini-SWE-agent by default (i.e., maximum step limit of 250 and maximum cost of \$3 per issue). We choose mini-SWE-agent to build on since it is not only simplistic (with $\sim$100 lines of code and only accessing bash commands) but also widely used. Unless otherwise stated, we use Claude 4.5 Sonnet (claude-sonnet-4-5-20250929(sonnet4.5) in our experiments and sample one patch per issue.
@@ 4.1 Main Results
SWE-bench Verified. Table 1 shows the results of Live-SWE-agent and prior agent approaches on SWE-bench Verified. We first observe that compared with mini-SWE-agent, across four different LLM backends, Live-SWE-agent consistently achieves a higher resolve rate. This demonstrates the improvement in performance enabled by allowing agents to create and use their own custom tools on the fly. Furthermore, Live-SWE-agent is able to achieve this with only a minimal increase in cost compared with the base mini-SWE-agent. In certain cases (e.g., in GPT-5), we even observed slight cost savings where the agent can improve the solving efficiency by replacing complex, multi-turn commands with customs tools that achieve the same functionality.
We also demonstrate the comparison between Live-SWE-agent and the best-performing agentic tools, including both state-of-the-art open-source solutions and proprietary commercial scaffolds, on SWE-bench Verified (Figure 1). We only report single-attempt results in Figure 1 without any test-time scaling to ensure a fair comparison. We observe that Live-SWE-agent using Gemini 3 Pro achieves a solve rate of 77.4% without test-time scaling, outperforming all existing agents on the SWE-bench Verified leaderboard [^1] https://www.swebench.com including state-of-the-art commercial solutions at the time of writing.
@@
SWE-Bench Pro. We also evaluate Live-SWE-agent on the public set of SWE-Bench Pro, containing 731 problems across 11 repositories and four programming languages (Python, Go, TypeScript, and JavaScript). Table 3 shows the performance of Live-SWE-agent compared to the state-of-the-art baseline of SWE-agent on SWE-Bench Pro deng2025swe. Different from mini-SWE-agent, which only has access to basic bash commands, SWE-agent provides handcrafted file viewing and editing tools and is the best performing approach on the SWE-Bench Pro. We choose Claude 4.5 Sonnet for SWE-agent because it ranks at the top of SWE-Bench Pro leaderboard [^2] https://scale.com/leaderboard/swe_bench_pro_public, as such, we also use it as the base LLM for Live-SWE-agent to enable a fair comparison. We observe that Live-SWE-agent is able to achieve better performance compared with SWE-agent, a specially designed agent implemented with close to 7,000 lines of code. We also compare with all other top-performing agents and LLMs in Figure 1 and observe that Live-SWE-agent is able to achieve the new state-of-the-art performance of 45.8% resolve rate on SWE-Bench Pro. This further demonstrates the superiority of the live scaffold design of Live-SWE-agent compared to existing agents that interact with a fixed set of manually-crafted tools.
@@ 4.2 Tools Analysis
In addition to the types of tools created, we also examine how the tools created can change across different repositories and languages. Figure 4(b) shows the visualization of the tools created in SWE-Bench Pro categorized based on repositories. In SWE-Bench Pro, there are 11 different repositories, and not surprisingly, there are common tools created across different repositories (i.e., the overlapping clusters in the figure). On the other hand, there are also specific tools for certain repositories, for example we see a distinct cluster (highlighted in orange diamonds) for openlibrary, a codebase for cataloging literature. Since openlibrary deals with lots of raw data of a specialized format, the custom tools generated by Live-SWE-agent are specifically tailored for that. We also observe similar results when we break down the tools based on programming language of the repositories in Figure 4(c). Different problems, environments, and languages may require completely different specialized tools, highlighting again how using the same set of tools for all problems is suboptimal.
@@ 4.3 Ablation
Different LLM backends. We now examine the effect of using different LLM backends for Live-SWE-agent. Table 5 compares the performance of the base mini-SWE-agent and Live-SWE-agent as we vary the LLM used on the same 50 problems used in the previous ablation experiment. We first observe that there are weaker LLMs for which using Live-SWE-agent to create custom tools on the fly even decreases performance. In particular, we see that for GPT-5-Nano, the results are significantly worse when using Live-SWE-agent compared to base mini-SWE-agent. After examining the trajectories, we found that GPT-5-Nano fails to understand the goal of creating custom tools and is often stuck in a loop thus leading to reduction in performance. This indicates that weaker LLMs may lack the high-level reasoning capabilities to synthesize useful tools on the fly. However, we see that Live-SWE-agent is able to improve more over the base mini-SWE-agent as we start to use more powerful models. Specifically, state-of-the-art LLMs such as Claude 4 Sonnet and GPT-5 achieves the highest relative resolve rate improvement when using Live-SWE-agent compared with mini-SWE-agent. This demonstrates the generalizability of Live-SWE-agent across high-performing LLMs and the potential for Live-SWE-agent to further improve performance especially as we continue to build more and more powerful LLMs.
@@ 4.4 Discussion and Future Work
Towards general, on the fly self-evolution. In this work, Live-SWE-agent primarily focuses on enabling agents to self-evolve through custom tool creation and usage. As discussed previously, the key idea of Live-SWE-agent, to improve and modify the agent itself on the fly, extends beyond creating new tools to also modifying the entire agent implementation. This includes the overall system prompt of the agent, how the agent interacts with the environment, and even the concrete workflow of the agent when it attempts to solve the issue. We believe that the entire agentic loop, much like any other software, can be modified and improved especially on the fly based on feedback and insights gained from solving a problem. Furthermore, we hope to extend the self-evolution loop across different tasks as well. Instead of discarding each evolved agent after a task is completed, we can save and serialize the useful tools and insights (via concepts like Skills skills) for future tasks. The agent can then load these useful tools and insights obtained while solving previous tasks on the fly to further improve performance and support continuous self-evolution across tasks.
@@ 5.1 Software Engineering Agents
Due to the huge design space for software agents, it can be extremely challenging and costly to build an optimal agent scaffold. As a result, a number of self-improving software agents have been proposed very recently, including the Self-Improving Coding Agent (SICA) (robeyns2025sica), Darwin-Gödel Machine (DGM) (zhang2025darwin), and Huxley-Gödel Machine (HGM) (wang2025huxley). However, such self-improving agents require costly offline training on known benchmarks, and may not be generalize well across different LLMs, benchmarks, and issue types. Beyond the software engineering domain, prior work llmastool; wang2024voyager; qiu2025alitageneralistagentenabling; qian2024creatortoolcreationdisentangling; wang2024troveinducingverifiableefficient explored using LLMs to create tools for general reasoning or embodied tasks, but they do not target real-world software engineering problems. In contrast, in this paper, we propose Live-SWE-agent, the first live software agent capable of performing practical self-evolution on-the-fly during runtime when solving real-world issues. In this way, Live-SWE-agent requires no offline training at all and can be easily generalized to different LLMs and issue domains. Moreover, it also demonstrated superior performance compared to all existing self-improving software agents.
@@ 5.2 Benchmarks for Software Engineering Agents
To evaluate and demonstrate the performance of software engineering agents, a large number of datasets have been proposed. SWE-bench (jimenez2023swe) is one of the earliest and most widely used benchmark datasets for software agents. Besides the initial SWE-bench dataset, which includes thousands of real-world GitHub issues, researchers have also built curated subsets with high-quality and representative issues to support faster and more reliable evaluation, including SWE-bench Lite (jimenez2023swe) and SWE-bench Verified (sbv). Since SWE-bench mostly focused on Python projects, researchers have proposed a number of SWE-bench-style benchmarks for projects in multiple languages for more comprehensive agent evaluation, including SWE-PolyBench (rashid2025swe), SWE-bench Multilingual (sbm), and Multi-SWE-bench (zan2025multi). Also, SWE-bench relies heavily on manual effort for collecting benchmark instances and setting up executable environments, and is hardly scalable; as a result, researchers have also leveraged software agents to streamline issue collection and environment setup to build live and scalable benchmarks, including SWE-bench-Live (zhang2025swe) and SWE-rebench badertdinov2025swe. More recently, Scale AI has also constructed SWE-Bench Pro (deng2025swe), which aims to capture more realistic, complex, enterprise-level issues than SWE-bench. For rigorous evaluation of Live-SWE-agent, our study involved multiple widely used benchmarks, including SWE-bench Verified, SWE-bench Multilingual, SWE-bench Pro.
@@ Appendix A Additional experimental settings
We now describe additional experimental settings used by Live-SWE-agent for evaluation. As discussed in Section 3, we implement Live-SWE-agent on top of the mini-SWE-agent framework. We use the same default hyperparameter of mini-SWE-agent at the time of writing (maximum step limit of 250 and maximum cost of \$3 per issue).
For Gemini 3 Pro, we use temperature of 1 since it is recommended by the developers [^4] https://ai.google.dev/gemini-api/docs/gemini-3. For any Anthropic models in our experiments, we follow mini-SWE-agent and use a temperature of 0.0. For any OpenAI models in our experiments (e.g., GPT-5, GPT-5-Mini, GPT-5-Nano), since temperature 0.0 sampling is not supported, we use temperature of 1. To the best of our knowledge, this is also the case used in mini-SWE-agent when they use OpenAI models.